OpenStudy (anonymous):

Solve the equation? √4x - 3 = x

OpenStudy (radar):

\[\sqrt{4x}=x+3\]

OpenStudy (radar):

square both sides.\[4x=x ^{2}+6x+9\]

OpenStudy (anonymous):

√4x = x +3 4x =(x+3)^2 4x=x^2+6x+9 x^2+2x+9=0 has no real solutions

OpenStudy (anonymous):

Last one .. Can someone answer it? √x - 6 + 4 = 9

OpenStudy (anonymous):

what's under the sqrt?

OpenStudy (anonymous):

only x?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

√x = 11

OpenStudy (anonymous):

x=121

OpenStudy (anonymous):

I checked Pre.CalSolved and it said 121 ... Wouldn't the sq root of that be 11?

OpenStudy (radar):

Consolidate terms\[x ^{2}+2x+9=0\] x= \[-2\pm \sqrt{4-4(9)}\over 2\]\[-2\pm \sqrt{-32}\over2\]\[-2\pm4\sqrt{-2}\over2\]\[-1\pm2\sqrt{2}i \]

OpenStudy (anonymous):

yes

OpenStudy (radar):

Provided the problem was as I copied